### v1

# Bertrand’s and Rodriguez Villegas’ Conjecture for real quartic Galois extensions of the rationals

Abstract.
The conjecture due to Bertrand and Rodriguez Villegas asserts that the 1-norm of the nonzero element in an exterior power of the units of a number field has a certain lower bound. For the exterior square case of totally real quartic extensions of the rationals, Costa and Friedman gave a lower bound of 0.802. We prove that the bound can be improved to 1.203 when the extension is further assumed to be Galois.
Key words and phrases:
Bertrand–Rodriguez Villegas conjecture, units
1. Introduction
While Dirichlet’s theorem determines the rank of the unit group in a given algebraic number field in terms of its Archimedean places, less is known about its geometry when viewed as a lattice in a Euclidean space via the associated logarithm embedding. The resulting lattice and its exterior powers are conjecturally subject to certain numerical constraints [1, 2]. We begin by introducing notation to clarify the nature of these numerical constraints.
Definition 1.1.
Let $L$ be a number field and $\mathcal{A}_{L}$ be the set of its Archimedean places. Then the logarithmic embedding of the units $\operatorname{LOG}:{\mathcal{O}_{L}^{*}}\rightarrow\mathbb{R}^{\mathcal{A}_{L}}$ into a Eucliedean space is defined by
$(\operatorname{LOG}(\gamma))_{v}:=e_{v}\log\lvert\gamma\rvert_{v}\text{\quad where\quad}e_{v}:=\begin{cases}1\ &\text{if}\ v\ \text{is real},\\ 2\ &\text{if}\ v\ \text{is complex}\end{cases}$
for $\gamma\in{\mathcal{O}_{L}^{*}}$ and $v\in\mathcal{A}_{L}$. Here, $\lvert\ \cdot\ \rvert_{v}$ is the absolute value associated to $v$ extending the absolute value on $\mathbb{Q}$.
Consider the orthonormal basis  $\{\delta^{v}\}_{v\in\mathcal{A}_{L}}$ of $\mathbb{R}^{\mathcal{A}_{L}}$ given by
$\delta_{w}^{v}:=\begin{cases}1\ &\text{if}\ \ w=v,\\ 0\ &\text{if}\ \ w\neq v\end{cases}$
for $w\in\mathcal{A}_{L}$. Let $\mathcal{A}_{L}^{[j]}$ be the set of subsets of $\mathcal{A}_{L}$ having cardinality $j$. For each  $I\in\mathcal{A}_{L}^{[j]}$, fix an ordering  $\{v_{1},v_{2},...,v_{j}\}$ of elements of $I$. Define
$\delta^{I}:=\delta^{v_{1}}\wedge\delta^{v_{2}}\wedge\cdots\wedge\delta^{v_{j}}$
to get an orthonormal basis  $\{\delta^{I}\}_{I\in\mathcal{A}_{L}^{[j]}}$ of  $\bigwedge^{j}\mathbb{R}^{\mathcal{A}_{L}}$. For $w=\sum\limits_{I\in\mathcal{A}_{L}^{[j]}}c_{I}\delta^{I}\in\bigwedge^{j}\mathbb{R}^{\mathcal{A}_{L}}$, define its 1-norm as
$\lVert w\rVert_{1}:=\sum\limits_{I\in\mathcal{A}_{L}^{[j]}}\lvert c_{I}\rvert.$
Our aim in this paper is to investigate the lower bound of  $\lVert w\rVert_{1}$ for nonzero element  $w$ of  $\bigwedge^{j}\operatorname{LOG}({\mathcal{O}_{L}^{*}})$. For a positive integer $n$, let $L_{n}$ denote the set of subfields of $\mathbb{R}$ that are Galois extension of $\mathbb{Q}$ of degree $n$. For such fields, $\operatorname{rank}({\mathcal{O}_{L}^{*}})=n-1$ by Dirichlet’s unit theorem. For $11\leq j\leq n-$ 1, define a real number $A_{n,j}$ as
(1.1) $A_{n,j}:=\inf_{\begin{subarray}{c}L\in L_{n}\\ w\in\bigwedge^{j}\operatorname{LOG}({\mathcal{O}_{L}^{*}})\backslash\{0\}\end{subarray}}\{\lVert w\rVert_{1}\}.$
This constant is directly related to the real Galois extension case of the Bertrand’s and Rodriguez Villegas’ conjecture. The conjecture stated in [2] is as follows.
Conjecture (Bertrand–Rodriguez Villegas).
There exists two absolute constants
 $c_{0}>0$ and $c_{1}>1$ such that for any number field $L$ and any  $j\in\mathbb{Z}_{>0}$, the following inequality holds
$\lVert w\rVert_{1}\geq c_{0}c_{1}^{j}$
for any nonzero $w\in\bigwedge^{j}\operatorname{LOG}({\mathcal{O}_{L}^{*}})\subset\bigwedge^{j}\mathbb{R}^{\mathcal{A}_{L}}$.
Only considering the number fields that are real Galois extensions of  $\mathbb{Q}$, a consequence of the conjecture is that
$A_{n,j}\geq c_{0}c_{1}^{j}$
holds, where the $c_{0},c_{1}$ are the constants satisfying the conjecture. For the specific case of  $n=4,j=2$, the work of Costa and Friedman [3] implies that
$A_{4,2}>2\sqrt{3}\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}\approx 0.802,$
as explained in § 2. To the best of the author’s knowledge, this was the largest known lower bound for $A_{4,2}$.
In this paper, we give a larger lower bound for $A_{4,2}$.
Theorem 1.2.
Let $L$ be a real quartic Galois extension of $\mathbb{Q}$. Then
$\lVert w\rVert_{1}>3\sqrt{3}\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}\approx 1.203$
for any nonzero $w\in\bigwedge^{2}\operatorname{LOG}({\mathcal{O}_{L}^{*}})$. In other words,
$A_{4,2}>3\sqrt{3}\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}\approx 1.203.$
As demonstrated in [2], Bertrand–Rodriguez Villegas conjecture is equivalent to Lehmer’s conjecture [6] when $j=1$, and to Zimmert’s theorem on regulators [8] when $j=\operatorname{rank}({\mathcal{O}_{L}^{*}})$. Hence the case of  $n=4,j=2$ is neither covered by Lehmer’s conjecture nor Zimmert’s theorem. While our result only covers this specific case, we obtained a better lower bound, which is $1.5$ times the previous known lower bound. In fact, for $L=\mathbb{Q}(\sqrt{2},\sqrt{5})$ and $w=\pm\operatorname{LOG}(\frac{1+\sqrt{5}}{2})\wedge\operatorname{LOG}(1+\sqrt{2})$, we have  $\lVert w\rVert_{1}=8\log(\frac{1+\sqrt{5}}{2})\log(1+\sqrt{2})$ and thus
$A_{4,2}\leq 8\log(\frac{1+\sqrt{5}}{2})\log(1+\sqrt{2})\approx 3.3930.$
Hence our lower bound is no less than a third of an upper bound of $A_{4,2}$.
The proof of the theorem is done separately for the case when  $\operatorname{Gal}(L/\mathbb{Q})$ is Klein four-group, and the case when  $\operatorname{Gal}(L/\mathbb{Q})$ is cyclic. For the Klein four-group case, we define a subgroup $E$ generated by the units of the quadratic subfields of $L$. Using the Galois module structure of $E$, we give a lower bound of  $\lVert w\rVert_{1}$ for nonzero element  $w$ of $\bigwedge^{2}\operatorname{LOG}(E)$ and extend this result to ${\mathcal{O}_{L}^{*}}$. For the cyclic group case, we use the structure of  ${\mathcal{O}_{L}^{*}}$ given by Hasse [5] to obtain a basis of $\bigwedge^{2}\operatorname{LOG}({\mathcal{O}_{L}^{*}})$. We then use the result of Pohst [7] to compute the lower bound.
In § 2 we derive the best lower bound of  $A_{4,2}$ known so far, from the work of Costa and Friedman [3]. In the remaining sections, we prove the main theorem to give the better lower bound for $A_{4,2}$. The case when  $\operatorname{Gal}(L/\mathbb{Q})$ is isomorphic to the Klein four-group is covered in § 3, and the case when  $\operatorname{Gal}(L/\mathbb{Q})$ is isomorphic to the cyclic group is covered in § 4.
Acknowledgement
The work of S.S. was supported by College of Natural Sciences Undergraduate Internship Program from Seoul National University. D.K. is supported by the National Research Foundation of Korea (NRF)111the grants No. RS-2023-00301976 and No. 2020R1C1C1A0100681913.
2. The lower bound due to Costa and Friedman
In this section, we derive the largest lower bound of  $A_{4,2}$ known so far, to the best of the author’s knowledge. Using Pohst’s result [7], Costa and Friedman [3] showed that for independent elements $\epsilon_{1},...,\epsilon_{j}$ of ${\mathcal{O}_{L}^{*}}$, the inequality
$\lVert\operatorname{LOG}(\epsilon_{1})\wedge\cdots\wedge\operatorname{LOG}(\epsilon_{j})\rVert_{2}>\left(\frac{n}{\gamma_{j}}\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}\right)^{j/2}$
holds for $1\leq j<n$, where $n=[L:\mathbb{Q}]$ and  $\gamma_{j}$ is Hermite’s constant in dimension $j$. If $j=n-2=\operatorname{rank}({\mathcal{O}_{L}^{*}})-1$, every element of  $\bigwedge^{n-2}\operatorname{LOG}({\mathcal{O}_{L}^{*}})$ can be written as  $d\cdot\operatorname{LOG}(\epsilon_{1})\wedge\cdots\wedge\operatorname{LOG}(\epsilon_{n-2})$ where $d\in\mathbb{Z}$ and $\operatorname{LOG}(\epsilon_{1}),...,\operatorname{LOG}(\epsilon_{n-1})$ are the basis of $\operatorname{LOG}({\mathcal{O}_{L}^{*}})$, as shown in the Lemma 28 of [2]. Therefore, for any nonzero element  $w$ of  $\bigwedge^{n-2}\operatorname{LOG}({\mathcal{O}_{L}^{*}})$, the inequality
$\lVert w\rVert_{1}\geq\lVert w\rVert_{2}>\left(\frac{n}{\gamma_{j}}\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}\right)^{j/2}$
holds. Restricting to the Galois extensions of $\mathbb{Q}$, we have
$A_{n,n-2}>\left(\frac{n}{\gamma_{n-2}}\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}\right)^{\frac{n-2}{2}}$
and in particular, since $\gamma_{2}=\frac{2}{\sqrt{3}}$, we obtain
$A_{4,2}>2\sqrt{3}\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}\approx 0.802.$
3. The Klein four-group Case
In this section, we prove the main theorem for the case when  $\operatorname{Gal}(L/\mathbb{Q})$ is isomorphic to the Klein four-group. Throughout this section, let $L$ be a real Galois extension of $\mathbb{Q}$ with $\operatorname{Gal}(L/\mathbb{Q})\cong\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$. Let $\mathbb{Q}(\sqrt{d_{1}}),\mathbb{Q}(\sqrt{d_{2}}),\mathbb{Q}(\sqrt{d_{3}})$ be three quadratic subfields of $L$, where $d_{1},d_{2},d_{3}$ are square-free integers. Then we can write $L=\mathbb{Q}(\sqrt{d_{1}},\sqrt{d_{2}})$. For $i=1,2,3$, let $\operatorname{Gal}(L/\mathbb{Q}(\sqrt{d_{i}}))=\{1,\sigma_{i}\}$. See Figure 1. Let  $u_{i}>1$ be the fundamental unit of $\mathbb{Q}(\sqrt{d_{i}})$. Without loss of generality, let $1<u_{1}<u_{2}<u_{3}$. If $i$ and $j$ are distinct integers in $\{1,2,3\}$, since $\sqrt{d_{j}}\not\in\mathbb{Q}(\sqrt{d_{i}})$$\sigma_{i}\in\operatorname{Gal}(L/\mathbb{Q}(\sqrt{d_{i}}))$ sends $\sqrt{d_{j}}$ to its conjugate $-\sqrt{d_{j}}$. Thus
(3.1) $u_{j}\sigma_{i}(u_{j})=\operatorname{N}_{\mathbb{Q}(\sqrt{d_{j}})/\mathbb{Q}}(u_{j})=\pm 1.$
$\mathbb{Q}$$\mathbb{Q}(\sqrt{d_{2}})$$\mathbb{Q}(\sqrt{d_{3}})$$\mathbb{Q}(\sqrt{d_{1}})$$L=\mathbb{Q}(\sqrt{d_{1}},\sqrt{d_{2}})$$\{id,\sigma_{1}\}$$\{id,\sigma_{2}\}$$\{id,\sigma_{3}\}$$\{id,\sigma_{1},\sigma_{2},\sigma_{3}\}$$\{id\}$
Figure 1. Subfields of non-cyclic extension $L$ and corresponding Galois groups
Lemma 3.1.
$u_{1},u_{2},u_{3}$ are multiplicatively independent. In other words, if integers $m_{1},m_{2},m_{3}$ satisfy $u_{1}^{m_{1}}u_{2}^{m_{2}}u_{3}^{m_{3}}=1$, then $m_{1}=m_{2}=m_{3}=0$.
Proof.
Assume that
$u_{1}^{m_{1}}u_{2}^{m_{2}}u_{3}^{m_{3}}=1$
for some $m_{1},m_{2},m_{3}\in\mathbb{Z}$. Then we have
$\sigma_{1}(u_{1})^{m_{1}}\sigma_{1}(u_{2})^{m_{2}}\sigma_{1}(u_{3})^{m_{3}}=1$
and thus
$(u_{1}\sigma_{1}(u_{1}))^{m_{1}}(u_{2}\sigma_{1}(u_{2}))^{m_{2}}(u_{3}\sigma_{1}(u_{3}))^{m_{3}}=1.$
By (3.1), $u_{2}\sigma_{1}(u_{2})=\pm 1$ and $u_{3}\sigma_{1}(u_{3})=\pm 1$. Therefore
$u_{1}^{2m_{1}}=\pm 1$
and thus $m_{1}=0$. Similarly, $m_{2}=m_{3}=0$. Therefore, $u_{1},u_{2},u_{3}$ are multiplicatively independent. ∎
Let $E=\{u_{1}^{m_{1}}u_{2}^{m_{2}}u_{3}^{m_{3}}|m_{1},m_{2},m_{3}\in\mathbb{Z}\}$, then we have the following lemma.
Lemma 3.2.
If $u\in{\mathcal{O}_{L}^{*}}$, then $u^{2}\in E$.
Proof.
From  $u\in{\mathcal{O}_{L}^{*}}$, we have  $\lvert N_{L/\mathbb{Q}}(u)\rvert=1$. The unit  $u$ of  $L$ and its conjugate  $\sigma_{1}(u)$ are units of  $L$, so  $u\sigma_{1}(u)$ is a unit of  $L$. Since  $u\sigma_{1}(u)=N_{L/\mathbb{Q}(\sqrt{d_{1}})}(u)\in\mathbb{Q}(\sqrt{d_{1}})$, it is also a unit of $\mathbb{Q}(\sqrt{d_{1}})$. Hence  $u\sigma_{1}(u)=\pm u_{1}^{m_{1}}$ for some $m_{1}\in\mathbb{Z}$. Similarly, $u\sigma_{2}(u)=\pm u_{2}^{m_{2}}$ and $u\sigma_{1}(u)=\pm u_{3}^{m_{3}}$ for some $m_{2},m_{3}\in\mathbb{Z}$. Then we have
$u^{2}=\lvert u^{2}\cdot N_{L/\mathbb{Q}}(u)\rvert=\lvert u^{3}\sigma_{1}(u)\sigma_{2}(u)\sigma_{3}(u)\rvert=\lvert\pm u_{1}^{m_{1}}u_{2}^{m_{2}}u_{3}^{m_{3}}\rvert=u_{1}^{m_{1}}u_{2}^{m_{2}}u_{3}^{m_{3}}$
and therefore $u^{2}\in E$. ∎
[…]
$\operatorname{LOG}(u_{1})=(\log(u_{1}),\ \ \log(u_{1}),-\log(u_{1}),-\log(u_{1}))$
$\operatorname{LOG}(u_{2})=(\log(u_{2}),-\log(u_{2}),\ \ \log(u_{2}),-\log(u_{2}))$
$\operatorname{LOG}(u_{3})=(\log(u_{3}),-\log(u_{3}),-\log(u_{3}),\ \ \log(u_{3})).$
Here, the order of the basis of $\mathbb{R}^{\mathcal{A}_{L}}$ is $\delta^{id},\delta^{\sigma_{1}},\delta^{\sigma_{2}},\delta^{\sigma_{3}}$. Then the wedge product of these $\operatorname{LOG}(u_{i})$’s are:
[…]
where $X_{1}=\log(u_{2})\log(u_{3}),X_{2}=\log(u_{1})\log(u_{3}),X_{3}=\log(u_{1})\log(u_{2})$ and the order of the basis of $\bigwedge^{j}\mathbb{R}^{\mathcal{A}_{L}}$ is $\delta^{id}\wedge\delta^{\sigma_{1}},\delta^{\sigma_{2}}\wedge\delta^{\sigma_{3}},\delta^{id}\wedge\delta^{\sigma_{3}},\delta^{\sigma_{1}}\wedge\delta^{\sigma_{2}},\delta^{id}\wedge\delta^{\sigma_{2}},\delta^{\sigma_{1}}\wedge\delta^{\sigma_{3}}$. Note that $X_{1}>X_{2}>X_{3}>0$.
For $w\in\bigwedge^{2}\operatorname{LOG}(E)$$w$ can be written as
$w=n_{1}\cdot\operatorname{LOG}(u_{2})\wedge\operatorname{LOG}(u_{3})+n_{2}\cdot\operatorname{LOG}(u_{1})\wedge\operatorname{LOG}(u_{3})+n_{3}\cdot\operatorname{LOG}(u_{1})\wedge\operatorname{LOG}(u_{2})$
where $n_{i}$’s are integers. Its 1-norm is
$\lVert w\rVert_{1}=$  $\,2(\lvert n_{2}X_{2}+n_{3}X_{3}\rvert+\lvert n_{2}X_{2}-n_{3}X_{3}\rvert+\lvert n_{1}X_{1}+n_{2}X_{2}\rvert+\lvert n_{1}X_{1}-n_{2}X_{2}\rvert$
$+\lvert n_{1}X_{1}+n_{3}X_{3}\rvert+\lvert n_{1}X_{1}-n_{3}X_{3}\rvert)$
$=$  $\,4(\max\{\lvert n_{2}X_{2}\rvert,\lvert n_{3}X_{3}\rvert\}+\max\{\lvert n_{1}X_{1}\rvert,\lvert n_{2}X_{2}\rvert\}+\max\{\lvert n_{1}X_{1}\rvert,\lvert n_{3}X_{3}\rvert\})$
where the second equality is from the following equation.
(3.2) $\lvert X+Y\rvert+\lvert X-Y\rvert=2\max\{\lvert X\rvert,\lvert Y\rvert\}\quad\forall X,Y\in\mathbb{R}.$
We have
$\max\{\lvert n_{1}X_{1}\rvert,\lvert n_{2}X_{2}\rvert\}+\max\{\lvert n_{1}X_{1}\rvert,\lvert n_{3}X_{3}\rvert\}\geq 2\lvert n_{1}X_{1}\rvert$
and thus
$4(\max\{\lvert n_{2}X_{2}\rvert,\lvert n_{3}X_{3}\rvert\}+\max\{\lvert n_{1}X_{1}\rvert,\lvert n_{2}X_{2}\rvert\}+\max\{\lvert n_{1}X_{1}\rvert,\lvert n_{3}X_{3}\rvert\})\geq 8\lvert n_{1}X_{1}\rvert$
holds. Similarly, the above inequality also holds when the right hand side is $8\lvert n_{2}X_{2}\rvert$ or $8\lvert n_{3}X_{3}\rvert$. Therefore
$\lVert w\rVert_{1}$ $=4(\max\{\lvert n_{2}X_{2}\rvert,\lvert n_{3}X_{3}\rvert\}+\max\{\lvert n_{1}X_{1}\rvert,\lvert n_{2}X_{2}\rvert\}+\max\{\lvert n_{1}X_{1}\rvert,\lvert n_{3}X_{3}\rvert\})$
$\geq 8\max\{\lvert n_{1}X_{1}\rvert,\lvert n_{2}X_{2}\rvert,\lvert n_{3}X_{3}\rvert\}$
and if $(n_{1},n_{2},n_{3})\neq 0$, then
(3.3) $\lVert w\rVert_{1}\geq 8\lvert X_{3}\rvert=8\log(u_{1})\log(u_{2}).$
For the case  $L=\mathbb{Q}(\sqrt{2},\sqrt{5})$, we have  $u_{1}=\frac{1+\sqrt{5}}{2}$,  $u_{2}=1+\sqrt{2}$, and  $u_{3}=3+\sqrt{10}$. A simple calculation shows that  $\sqrt{u_{1}u_{2}u_{3}}$ is in  ${\mathcal{O}_{L}^{*}}$, and it is easy to verify that  $u_{1},u_{2},\sqrt{u_{1}u_{2}u_{3}}$ generate  ${\mathcal{O}_{L}^{*}}$. Then the group  $\bigwedge^{2}\operatorname{LOG}({\mathcal{O}_{L}^{*}})$ is generated by $\operatorname{LOG}(u_{1})\wedge\operatorname{LOG}(u_{2})$$\frac{1}{2}(\operatorname{LOG}(u_{1})\wedge\operatorname{LOG}(u_{2})+\operatorname{LOG}(u_{1})\wedge\operatorname{LOG}(u_{3}))$ and $\frac{1}{2}(\operatorname{LOG}(u_{1})\wedge\operatorname{LOG}(u_{2})+\operatorname{LOG}(u_{2})\wedge\operatorname{LOG}(u_{3}))$. Thus for any element $w$ of $\bigwedge^{2}\operatorname{LOG}({\mathcal{O}_{L}^{*}})$$2w\in\bigwedge^{2}\operatorname{LOG}(E)$. Hence we have
(3.4) $\lVert w\rVert_{1}=\frac{1}{2}\lVert 2w\rVert_{1}\geq 4\log\left(\frac{1+\sqrt{5}}{2}\right)\log(1+\sqrt{2})\approx 1.697.$
For the case $L=\mathbb{Q}(\sqrt{5},\sqrt{13})$$u_{1}=\frac{1+\sqrt{5}}{2},u_{2}=\frac{3+\sqrt{13}}{2},u_{3}=8+\sqrt{65}$. Similarly, $u_{1},u_{2},\sqrt{u_{1}u_{2}u_{3}}$ generate  ${\mathcal{O}_{L}^{*}}$ and therefore
(3.5) $\lVert w\rVert_{1}=\frac{1}{2}\lVert 2w\rVert_{1}\geq 4\log\left(\frac{1+\sqrt{5}}{2}\right)\log(\frac{3+\sqrt{13}}{2})\approx 2.300$
for any nonzero element $w$ of $\bigwedge^{2}\operatorname{LOG}({\mathcal{O}_{L}^{*}})$.
Now consider the case where $L\neq\mathbb{Q}(\sqrt{2},\sqrt{5})$ and $L\neq\mathbb{Q}(\sqrt{5},\sqrt{13})$. By Lemma 3.2, for any $u\in{\mathcal{O}_{L}^{*}}$$u^{2}\in E$ and thus $2\operatorname{LOG}(u)\in\operatorname{LOG}(E)$. Hence if nonzero $w$ is in $\bigwedge^{2}\operatorname{LOG}({\mathcal{O}_{L}^{*}})$, then $4w\in\bigwedge^{2}\operatorname{LOG}(E)$. We have
(3.6) $\lVert w\rVert_{1}=\frac{1}{4}\lvert\lvert 4w\rvert\rvert_{1}\geq 2\log(u_{1})\log(u_{2})$
by (3.3). Excluding $\mathbb{Q}(\sqrt{2},\sqrt{5})$ and $\mathbb{Q}(\sqrt{5},\sqrt{13})$, $\mathbb{Q}(\sqrt{3},\sqrt{5})$ has the smallest possible value of  $\log(u_{1})\log(u_{2})$, which is  $\log\left(\frac{1+\sqrt{5}}{2}\right)\log(2+\sqrt{3})$. This is from the Lemma 3.3 below and the fact that $\log\left(\frac{1+\sqrt{5}}{2}\right)\log(2+\sqrt{3})<\log(1+\sqrt{2})\log(\frac{3+\sqrt{13}}{2})$. Hence from (3.6), we see that
$\lVert w\rVert_{1}\geq 2\log\left(\frac{1+\sqrt{5}}{2}\right)\log(2+\sqrt{3})\approx 1.267.$
With (3.4) and (3.5), we conclude that if $\operatorname{Gal}(L/\mathbb{Q})$ is Klein four-group,
$\lVert w\rVert_{1}\geq 2\log\left(\frac{1+\sqrt{5}}{2}\right)\log(2+\sqrt{3})>3\sqrt{3}\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}$
for any nonzero $w\in\bigwedge^{2}\operatorname{LOG}({\mathcal{O}_{L}^{*}})$.
Lemma 3.3.
For square-free integer $m$, let $v_{m}>1$ be the fundamental unit of $\mathbb{Q}(\sqrt{m})$. Then  $\frac{1+\sqrt{5}}{2},1+\sqrt{2},\frac{3+\sqrt{13}}{2},2+\sqrt{3}$ are the four smallest possible values of $v_{m}$.
Proof.
First, consider the case $m\equiv 2,3\pmod{4}$. The fundamental unit $v_{m}$ is of the form $a+b\sqrt{m}$ where  $a,b$ are the smallest positive integers satisfying $a^{2}-mb^{2}=\pm 1$. If $m\geq 6$, then
$v_{m}\geq\sqrt{mb^{2}-1}+b\sqrt{m}\geq\sqrt{m-1}+\sqrt{m}\geq\sqrt{5}+\sqrt{6}>2+\sqrt{3}.$
Now consider the case $m\equiv 1\pmod{4}$. Then $v_{m}$ is of the form $\frac{a+b\sqrt{m}}{2}$ where  $a,b$ are the smallest positive integers satisfying $a^{2}-mb^{2}=\pm 4$. If $m\geq 17$,
$v_{m}\geq\frac{\sqrt{mb^{2}-4}+b\sqrt{m}}{2}\geq\frac{\sqrt{m-4}+\sqrt{m}}{2}\geq\frac{\sqrt{13}+\sqrt{17}}{2}>2+\sqrt{3}.$
Hence $v_{m}\leq 2+\sqrt{3}$ only for $m=2,3,5,13$. For those $m$, we have  $v_{5}<v_{2}<v_{13}<v_{3}$. ∎
4. The Cyclic Group Case
In this section, we prove the main theorem for the case when  $\operatorname{Gal}(L/\mathbb{Q})$ is isomorphic to the cyclic group. Throughout this section, let $L$ be a real Galois extension of $\mathbb{Q}$ with $\operatorname{Gal}(L/\mathbb{Q})\cong\mathbb{Z}/4\mathbb{Z}$. Denote the Galois group as  $G$. Let $\operatorname{Gal}(L/\mathbb{Q})$ be generated by $\sigma$. Let $l$ be the unique quadratic subfield of $L$. See Figure 2. Let $u_{l}>1$ be the fundamental unit of $l$. Then ${\mathcal{O}_{l}^{*}}=\{\pm u_{l}^{m}\ |\ m\in\mathbb{Z}\}$. First we define relative units as in [5].
$\mathbb{Q}$$l$$L$$\{1,\sigma^{2}\}$$G=\{1,\sigma,\sigma^{2},\sigma^{3}\}$$\{1\}$
Figure 2. Subfields of cyclic extension $L$ and corresponding Galois groups
Definition 4.1.
If $w\in{\mathcal{O}_{L}^{*}}$ satisfies $N_{L/l}(w)=\pm 1$, we say that $w$ is a relative unit of $L$.
Let $E_{L/l}$ be a group consisting of all relative units of $L$. As in [4], let $E^{L}$ be a sub $G$-module of ${\mathcal{O}_{L}^{*}}$ generated by elements of  ${\mathcal{O}_{l}^{*}}$ and $E_{L/l}$. Let $Q=[{\mathcal{O}_{L}^{*}}:E^{L}]$. The following lemmas are from Hasse [5].
Lemma 4.2.
The index $Q$ is either 1 or 2. Furthermore,  $Q=2$ if and only if there exists $u_{*}\in{\mathcal{O}_{L}^{*}}$ such that $N_{L/l}(u_{*})=\pm u_{l}$.
Proof.
See Satz 16, p.37 of Hasse [5]. ∎
Lemma 4.3.
There exists an element $u_{0}$ of $E_{L/l}$ such that $-1,u_{0}$ and its conjugate $\sigma(u_{0})$ generates $E_{L/l}$.
Proof.
See Satz 22, p.40 of Hasse [5]. ∎
Define $u_{0}$ as in the lemma. We have
$E^{L}=\{\pm u_{l}^{m_{1}}u_{0}^{m_{2}}\sigma(u_{0})^{m_{3}}\ |\ (m_{1},m_{2},m_{3})\in\mathbb{Z}^{3}\}.$
We first compute the 1-norm of the elements of $\bigwedge^{2}\operatorname{LOG}(E^{L})$. Note that since $u_{0}$ is relative unit of $L$, $\sigma^{2}(u_{0})=\pm\frac{1}{u_{0}}$. We can compute $\operatorname{LOG}$’s as follows:
(4.1) $\begin{split}\operatorname{LOG}(u_{l})=(W_{1},-W_{1},W_{1},-W_{1})\\ \operatorname{LOG}(u_{0})=(W_{2},W_{3},-W_{2},-W_{3})\\ \operatorname{LOG}(\sigma(u_{0}))=(W_{3},-W_{2},-W_{3},W_{2})\end{split}$
where $W_{1}=\log(u_{l}),W_{2}=\log(\lvert u_{0}\rvert),W_{3}=\log(\lvert\sigma(u_{0})\rvert)$. Here, the order of the basis of $\mathbb{R}^{\mathcal{A}_{L}}$ is $\delta^{id},\delta^{\sigma},\delta^{\sigma^{2}},\delta^{\sigma^{3}}$. Note that $W_{1},W_{2},W_{3}$ are nonzero since $u_{l},u_{0},\sigma(u_{0})$ are not ±$\pm 1$. Then the wedge product of these terms are:
$\operatorname{LOG}(u_{l})\wedge\operatorname{LOG}(u_{0})=$ $(\ \ Y_{4},-Y_{4},Y_{5},\ \ Y_{5},-Y_{2},\ \ Y_{3})$
$\operatorname{LOG}(u_{l})\wedge\operatorname{LOG}(\sigma(u_{0}))=$ $(-Y_{5},\ \ Y_{5},Y_{4},\ \ Y_{4},-Y_{3},-Y_{2})$
[…]
where $Y_{1}=W_{2}^{2}+W_{3}^{2},Y_{2}=2W_{1}W_{2},Y_{3}=2W_{1}W_{3},Y_{4}=W_{1}W_{2}+W_{1}W_{3},Y_{5}=W_{1}W_{2}-W_{1}W_{3}$. Here the order of the basis of $\bigwedge^{2}\mathbb{R}^{\mathcal{A}_{L}}$ is $\delta^{id}\wedge\delta^{\sigma},\delta^{\sigma^{2}}\wedge\delta^{\sigma^{3}},\delta^{id}\wedge\delta^{\sigma^{3}},\delta^{\sigma}\wedge\delta^{\sigma^{2}},\delta^{id}\wedge\delta^{\sigma^{2}},\delta^{\sigma}\wedge\delta^{\sigma^{3}}$.
For any nonzero element  $w\in\bigwedge^{2}\operatorname{LOG}(E^{L})$, it can be written as
$w=$ $\,n_{1}\cdot\operatorname{LOG}(u_{l})\wedge\operatorname{LOG}(u_{0})+n_{2}\cdot\operatorname{LOG}(u_{l})\wedge\operatorname{LOG}(\sigma(u_{0}))$
$+n_{3}\cdot\operatorname{LOG}(u_{0})\wedge\operatorname{LOG}(\sigma(u_{0}))$
where $n_{1},n_{2},n_{3}$ are integers. Then $\lVert w\rVert_{1}=f(n_{1},n_{2},n_{3})$ where the map $f:\mathbb{Z}^{3}\rightarrow\mathbb{R}$ is given as
$f(n_{1},n_{2},n_{3})=$ $\,\lvert n_{1}Y_{4}-n_{2}Y_{5}-n_{3}Y_{1}\rvert+\lvert-n_{1}Y_{4}+n_{2}Y_{5}-n_{3}Y_{1}\rvert$
$+\lvert n_{1}Y_{5}+n_{2}Y_{4}+n_{3}Y_{1}\rvert+\lvert n_{1}Y_{5}+n_{2}Y_{4}-n_{3}Y_{1}\rvert$
$+\lvert-n_{1}Y_{2}-n_{2}Y_{3}\rvert+\lvert n_{1}Y_{3}-n_{2}Y_{2}\rvert.$
Applying (3.2), above equation becomes
$f(n_{1},n_{2},n_{3})=$ $2\max\{\lvert n_{1}Y_{4}-n_{2}Y_{5}\rvert,\lvert n_{3}Y_{1}\rvert\}+2\max\{\lvert n_{1}Y_{5}+n_{2}Y_{4}\rvert,\lvert n_{3}Y_{1}\rvert\}$
$+\lvert-n_{1}Y_{2}-n_{2}Y_{3}\rvert+\lvert n_{1}Y_{3}-n_{2}Y_{2}\rvert.$
In particular, we have
(4.2) $f(n_{1},n_{2},n_{3})\geq 4\lvert n_{3}Y_{3}\rvert=4\lvert n_{3}\rvert(W_{2}^{2}+W_{3}^{2})$
and
$f(n_{1},n_{2},n_{3})\geq$  $\,2\lvert n_{1}Y_{4}-n_{2}Y_{5}\rvert+2\lvert n_{1}Y_{5}+n_{2}Y_{4}\rvert$
$+\lvert-n_{1}Y_{2}-n_{2}Y_{3}\rvert+\lvert n_{1}Y_{3}-n_{2}Y_{2}\rvert.$
If $(n_{1},n_{2})\neq(0,0)$, then by Lemma 4.4 below,
$f(n_{1},n_{2},n_{3})$ $\geq 2\lvert Y_{4}\rvert+2\lvert Y_{5}\rvert+\lvert Y_{2}\rvert+\lvert Y_{3}\rvert$
$=2\lvert W_{1}W_{2}+W_{1}W_{3}\rvert+2\lvert W_{1}W_{2}-W_{1}W_{3}\rvert+2\lvert W_{1}W_{2}\rvert+2\lvert W_{1}W_{3}\rvert$
and applying (3.2) once more, we have
(4.3) $f(n_{1},n_{2},n_{3})\geq 2W_{1}(2\max\{\lvert W_{2}\rvert,\lvert W_{3}\rvert\}+\lvert W_{2}\rvert+\lvert W_{3}\rvert)$
Using these results, we prove the Theorem 1.2 for each case $Q=1$ and $Q=2$.
Lemma 4.4.
Let $X,Y\in\mathbb{R}$. For any $(m,n)\in\mathbb{Z}^{2}\backslash\{(0,0)\}$, the following inequality
$\lvert mX+nY\rvert+\lvert nX-mY\rvert\geq\lvert X\rvert+\lvert Y\rvert$
holds.
Proof.
The inequality is trivial if $X=0$ or $Y=0$. Assume that they are both nonzero. We may further assume that they are positive and $X\geq Y$. Let $Z:=Y/X\leq 1$. Since
$\lvert mX+nY\rvert+\lvert nX-mY\rvert=\lvert X\rvert(\lvert m+nZ\rvert+\lvert n-mZ\rvert),$
we only need to show the inequality
$\lvert m+nZ\rvert+\lvert n-mZ\rvert\geq 1+Z.$
Without loss of generality, assume $m\geq 0$. If $m=0$$\lvert nZ\rvert+\lvert n\rvert\geq 1+Z$ for any nonzero integer $n$. If $m\geq 2$,
$\lvert m+nZ\rvert+\lvert n-mZ\rvert$ $\geq\lvert m+nZ\rvert+\lvert Z(n-mZ)\rvert$
$\geq\lvert(m+nZ)-Z(n-mZ)\rvert$
$=\lvert m(1+Z^{2})\rvert$
$\geq m$
$\geq 2$
$\geq 1+Z.$
Now consider the case $m=1$. The case $n=0$ is trivial. If $n\leq-1$, then
$\lvert 1+nZ\rvert+\lvert n-Z\rvert\geq{n-Z}\geq 1+Z$
and if $n\geq 1$,
$\lvert 1+nZ\rvert+\lvert n-Z\rvert\geq{1+nZ}\geq 1+Z.$
4.1. The Case Q=1
Since ${\mathcal{O}_{L}^{*}}=E^{L}$, for any nonzero $w\in\bigwedge^{2}\operatorname{LOG}({\mathcal{O}_{L}^{*}})$, its norm is of the form  $\lVert w\rVert_{1}=f(n_{1},n_{2},n_{3})$ where $(n_{1},n_{2},n_{3})\neq(0,0,0)$. If $n_{3}\neq 0$, then (4.2) gives
$\lVert w\rVert_{1}\geq 4(W_{2}^{2}+W_{3}^{2})$
and if $n_{3}=0$, then $(n_{1},n_{2})\neq(0,0)$, so (4.3) gives
$\lVert w\rVert_{1}\geq 2W_{1}(2\max\{\lvert W_{2}\rvert,\lvert W_{3}\rvert\}+\lvert W_{2}\rvert+\lvert W_{3}\rvert).$
Therefore we have
(4.4) $\lVert w\rVert_{1}\geq\min\{4(W_{2}^{2}+W_{3}^{2}),2W_{1}(2\max\{\lvert W_{2}\rvert,\lvert W_{3}\rvert\}+\lvert W_{2}\rvert+\lvert W_{3}\rvert)\}$
for any nonzero $w\in\bigwedge\operatorname{LOG}({\mathcal{O}_{L}^{*}})$.
We use the following theorem from Pohst [7] to give a lower bound for (4.4).
Theorem 4.5.
Let $L$ be totally real number field and $u\in{\mathcal{O}_{L}^{*}}$ with $u\neq\pm 1$. Then
$\lVert\operatorname{LOG}(u)\rVert_{2}^{2}\geq[L:\mathbb{Q}]\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}$
holds.
Proof.
Bemerkung in p.102 of Pohst [7] states that the Lemma in p.98 applies for  $M$ of the eqaution (10) in p.97. Then the inequality (11) of the Lemma gives the required result. ∎
Then from the above theorem and (4.1),
(4.5) $W_{2}^{2}+W_{3}^{2}=\frac{1}{2}\lVert\operatorname{LOG}(u_{0})\rVert_{2}^{2}\geq 2\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}.$
Under this constraint and the fact that $W_{2},W_{3}$ are nonzero, an elementary consideration yields
$2\max\{\lvert W_{2}\rvert,\lvert W_{3}\rvert\}+\lvert W_{2}\rvert+\lvert W_{3}\rvert>3\sqrt{2}\log\left(\frac{1+\sqrt{5}}{2}\right).$
Since $u_{l}$ is a fundamental unit of quadratic field, by Lemma 3.3,
(4.6) $W_{1}=\log(u_{l})\geq\log\left(\frac{1+\sqrt{5}}{2}\right)$
holds, and with previous inequality,
$2W_{1}(2\max\{\lvert W_{2}\rvert,\lvert W_{3}\rvert\}+\lvert W_{2}\rvert+\lvert W_{3}\rvert)>6\sqrt{2}\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}.$
With (4.5), the inequality (4.4) becomes
(4.7) $\lVert w\rVert_{1}\geq\min\left\{8\log\left(\frac{1+\sqrt{5}}{2}\right)^{2},6\sqrt{2}\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}\right\}=8\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}$
for any nonzero $w\in\bigwedge^{2}\operatorname{LOG}({\mathcal{O}_{L}^{*}})$. Thus Theorem 1.2 holds for the case $Q=1$.
4.2. The Case Q=2
We use the following lemma from Hasse [5].
Lemma 4.6.
If $Q=2$, there exists a unique unit $u_{*}>0$ of $L$ that satisfy
(4.8) $N_{L/l}(u_{*})=u_{*}\sigma^{2}(u_{*})=\pm u_{l},\quad u_{*}\sigma(u_{*})=\pm u_{0}.$
Proof.
See Satz 28, p.45 of Hasse [5]. ∎
Define $u_{*}$ as in the lemma. Then
(4.9) $u_{*}^{2}=\frac{u_{*}\sigma^{2}(u_{*})\cdot u_{*}\sigma(u_{*})}{\sigma(u_{*}\sigma(u_{*}))}=\pm\frac{u_{l}u_{0}}{\sigma(u_{0})}$
and therefore  $\sigma(u_{0})=\pm u_{l}u_{0}u_{*}^{-2}$. The proof of Lemma 4.2 shows that when $Q=2$, every unit $u\in{\mathcal{O}_{L}^{*}}$ can be written uniquely as the form ±$\pm u_{*}^{n_{0}}u_{l}^{n_{1}}u_{0}^{n_{2}}\sigma(u_{0})^{n_{3}}$ where $n_{0}$ is 00 or $1$ and $n_{1},n_{2},n_{3}$ are integers. Thus  ${\mathcal{O}_{L}^{*}}$ is generated by $-1,u_{l},u_{0},u_{*}$. Hence  $\bigwedge^{2}\operatorname{LOG}({\mathcal{O}_{L}^{*}})$ is generated by $\operatorname{LOG}(u_{l})\wedge\operatorname{LOG}(u_{0}),\operatorname{LOG}(u_{l})\wedge\operatorname{LOG}(u_{*})$ and  $\operatorname{LOG}(u_{0})\wedge\operatorname{LOG}(u_{*})$. From (4.9), we also have
(4.10) $\operatorname{LOG}(u_{*})=\frac{1}{2}(\operatorname{LOG}(u_{l})+\operatorname{LOG}(u_{0})-\operatorname{LOG}(\sigma(u_{0})))$
and thus
$\operatorname{LOG}(u_{l})\wedge\operatorname{LOG}(u_{*})$ $=\frac{1}{2}(\operatorname{LOG}(u_{l})\wedge\operatorname{LOG}(u_{0})-\operatorname{LOG}(u_{l})\wedge\operatorname{LOG}(\sigma(u_{0}))),$
$\operatorname{LOG}(u_{0})\wedge\operatorname{LOG}(u_{*})$ $=-\frac{1}{2}(\operatorname{LOG}(u_{l})\wedge\operatorname{LOG}(u_{0})+\operatorname{LOG}(u_{0})\wedge\operatorname{LOG}(\sigma(u_{0}))).$
Hence for any $w\in\bigwedge^{2}\operatorname{LOG}({\mathcal{O}_{L}^{*}})$,
$2w$ $=n_{1}\cdot\operatorname{LOG}(u_{l})\wedge\operatorname{LOG}(u_{0})+n_{2}\cdot\operatorname{LOG}(u_{l})\wedge\operatorname{LOG}(\sigma(u_{0}))$
$+n_{3}\cdot\operatorname{LOG}(u_{0})\wedge\operatorname{LOG}(\sigma(u_{0}))$
where $n_{1},n_{2},n_{3}$ are integers and $n_{1}+n_{2}+n_{3}$ is even and thus
(4.11) $\lVert w\rVert_{1}=\frac{1}{2}f(n_{1},n_{2},n_{3}).$
We now proceed to give a lower bound for $f(n_{1},n_{2},n_{3})$ where $n_{1}+n_{2}+n_{3}$ is even and $(n_{1},n_{2},n_{3})\neq(0,0,0)$. If $(n_{1},n_{2})=0$, then $\lvert n_{3}\rvert\geq 2$, so (4.2) gives
$f(n_{1},n_{2},n_{3})\geq 8(W_{2}^{2}+W_{3}^{2}).$
Since (4.5) still holds in the $Q=2$ case, we have
(4.12) $f(n_{1},n_{2},n_{3})\geq 8\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}.$
Note that (4.6) also still holds.
Now consider the case $(n_{1},n_{2})\neq(0,0)$. Since $\operatorname{LOG}(u_{l}),\operatorname{LOG}(u_{0}),\operatorname{LOG}(\sigma(u_{0}))~{}$ are perpendicular as seen in (4.1), we have
$\lVert\operatorname{LOG}(u_{*})\rVert_{2}^{2}$ $=\frac{1}{4}(\lVert\operatorname{LOG}(u_{l})\rVert_{2}^{2}+\lVert\operatorname{LOG}(u_{0})\rVert_{2}^{2}+\lVert\operatorname{LOG}(\sigma(u_{0}))\rVert_{2}^{2})$
$=W_{1}^{2}+W_{2}^{2}+W_{3}^{2}$
from (4.10). Then Theorem 4.5 gives
$W_{1}^{2}+W_{2}^{2}+W_{3}^{2}\geq 4\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}.$
Under the constraint of above inequality, (4.5), (4.6), and the fact that  $W_{1},W_{2},W_{3}$ are nonzero, an elementary consideration yields
$2W_{1}(2\max\{\lvert W_{2}\rvert,\lvert W_{3}\rvert\}+\lvert W_{2}\rvert+\lvert W_{3}\rvert)>6\sqrt{3}\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}.$
With (4.3), we have
$f(n_{1},n_{2},n_{3})>6\sqrt{3}\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}.$
Since $8\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}>6\sqrt{3}\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}$, by (4.12), the above inequality holds for any $(n_{1},n_{2},n_{3})\neq(0,0,0)$ with $n_{1}+n_{2}+n_{3}$ even. Therefore, with (4.11),
$\lVert w\rVert_{1}>3\sqrt{3}\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}$
for any nonzero $w\in\bigwedge^{2}\operatorname{LOG}({\mathcal{O}_{L}^{*}})$. This concludes the proof of Theorem 1.2.

### v2

@@
 For the exterior square case of totally real quartic extensions of the rationals, Costa and Friedman gave a lower bound of 0.802.
-We prove that the bound can be improved to 1.203 when the extension is further assumed to be Galois.
+We prove that the bound can be improved to 1.134 when the extension is further assumed to be Galois.
 Key words and phrases:
-Bertrand–Rodriguez Villegas conjecture, units
+Bertrand–Rodriguez Villegas conjecture, units, heights, Lehmer’s conjecture
 1. Introduction
@@
 The resulting lattice and its exterior powers are conjecturally subject to certain numerical constraints [1, 2].
-We begin by introducing notation to clarify the nature of these numerical constraints.
+We begin by introducing notation to clarify the nature of this numerical constraint.
 Definition 1.1.
@@
 to get an orthonormal basis  $\{\delta^{I}\}_{I\in\mathcal{A}_{L}^{[j]}}$ of  $\bigwedge^{j}\mathbb{R}^{\mathcal{A}_{L}}$.
-For $w=\sum\limits_{I\in\mathcal{A}_{L}^{[j]}}c_{I}\delta^{I}\in\bigwedge^{j}\mathbb{R}^{\mathcal{A}_{L}}$, define its 1-norm as
-$\lVert w\rVert_{1}:=\sum\limits_{I\in\mathcal{A}_{L}^{[j]}}\lvert c_{I}\rvert.$
-Our aim in this paper is to investigate the lower bound of  $\lVert w\rVert_{1}$ for nonzero element  $w$ of  $\bigwedge^{j}\operatorname{LOG}({\mathcal{O}_{L}^{*}})$.
+For $w=\sum\limits_{I\in\mathcal{A}_{L}^{[j]}}c_{I}\delta^{I}\in\bigwedge^{j}\mathbb{R}^{\mathcal{A}_{L}}$, define its 1-norm and 2-norm as
+$\lVert w\rVert_{1}:=\sum\limits_{I\in\mathcal{A}_{L}^{[j]}}\lvert c_{I}\rvert,\qquad\lVert w\rVert_{2}:=\bigg(\sum\limits_{I\in\mathcal{A}_{L}^{[j]}}c_{I}^{2}\bigg)^{1/2}$
+respectively.
+Our aim in this paper is to investigate a lower bound for  $\lVert w\rVert_{1}$ for a nonzero element  $w$ of  $\bigwedge^{j}\operatorname{LOG}({\mathcal{O}_{L}^{*}})$.
 For a positive integer $n$, let $L_{n}$ denote the set of subfields of $\mathbb{R}$ that are Galois extension of $\mathbb{Q}$ of degree $n$.
 For such fields, $\operatorname{rank}({\mathcal{O}_{L}^{*}})=n-1$ by Dirichlet’s unit theorem.
-For $11\leq j\leq n-$ 1, define a real number $A_{n,j}$ as
+For $1\leq j\leq n-1$, define a real number $A_{n,j}$ as
 (1.1) $A_{n,j}:=\inf_{\begin{subarray}{c}L\in L_{n}\\ w\in\bigwedge^{j}\operatorname{LOG}({\mathcal{O}_{L}^{*}})\backslash\{0\}\end{subarray}}\{\lVert w\rVert_{1}\}.$
-This constant is directly related to the real Galois extension case of the Bertrand’s and Rodriguez Villegas’ conjecture. The conjecture stated in [2] is as follows.
+This constant is directly related to the real Galois extension case of Bertrand’s and Rodriguez Villegas’ conjecture. The conjecture stated in [2] is as follows.
 Conjecture (Bertrand–Rodriguez Villegas).
-There exists two absolute constants
+There exist two absolute constants
  $c_{0}>0$ and $c_{1}>1$ such that for any number field $L$ and any  $j\in\mathbb{Z}_{>0}$, the following inequality holds
@@
 holds, where the $c_{0},c_{1}$ are the constants satisfying the conjecture. For the specific case of  $n=4,j=2$, the work of Costa and Friedman [3] implies that
-$A_{4,2}>2\sqrt{3}\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}\approx 0.802,$
-as explained in § 2. To the best of the author’s knowledge, this was the largest known lower bound for $A_{4,2}$.
+$A_{4,2}>2\sqrt{3}\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right)\approx 0.802,$
+as explained in § 2. Throughout this paper, we write $\log^{2}(x)$ to mean $(\log(x))^{2}$. To the best of the authors’ knowledge, this was the largest known lower bound for $A_{4,2}$.
 In this paper, we give a larger lower bound for $A_{4,2}$.
@@
 Let $L$ be a real quartic Galois extension of $\mathbb{Q}$. Then
-$\lVert w\rVert_{1}>3\sqrt{3}\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}\approx 1.203$
+$\lVert w\rVert_{1}\geq 2\sqrt{6}\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right)\approx 1.134$
 for any nonzero $w\in\bigwedge^{2}\operatorname{LOG}({\mathcal{O}_{L}^{*}})$. In other words,
-$A_{4,2}>3\sqrt{3}\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}\approx 1.203.$
-As demonstrated in [2], Bertrand–Rodriguez Villegas conjecture is equivalent to Lehmer’s conjecture [6] when $j=1$, and to Zimmert’s theorem on regulators [8] when $j=\operatorname{rank}({\mathcal{O}_{L}^{*}})$.
+$A_{4,2}\geq 2\sqrt{6}\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right)\approx 1.134.$
+As demonstrated in [2], the Bertrand–Rodriguez Villegas conjecture is equivalent to Lehmer’s conjecture [7] when $j=1$, and to Zimmert’s theorem on regulators [9] when $j=\operatorname{rank}({\mathcal{O}_{L}^{*}})$.
 Hence the case of  $n=4,j=2$ is neither covered by Lehmer’s conjecture nor Zimmert’s theorem.
-While our result only covers this specific case, we obtained a better lower bound, which is $1.5$ times the previous known lower bound.
-In fact, for $L=\mathbb{Q}(\sqrt{2},\sqrt{5})$ and $w=\pm\operatorname{LOG}(\frac{1+\sqrt{5}}{2})\wedge\operatorname{LOG}(1+\sqrt{2})$, we have  $\lVert w\rVert_{1}=8\log(\frac{1+\sqrt{5}}{2})\log(1+\sqrt{2})$ and thus
-$A_{4,2}\leq 8\log(\frac{1+\sqrt{5}}{2})\log(1+\sqrt{2})\approx 3.3930.$
-Hence our lower bound is no less than a third of an upper bound of $A_{4,2}$.
-The proof of the theorem is done separately for the case when  $\operatorname{Gal}(L/\mathbb{Q})$ is Klein four-group, and the case when  $\operatorname{Gal}(L/\mathbb{Q})$ is cyclic.
+While our result only covers this specific case, we obtained a better lower bound, which is $\sqrt{2}$ times the previous known lower bound.
+In fact, for $L=\mathbb{Q}(\sqrt{2},\sqrt{5})$ and $w=\pm\operatorname{LOG}(\tfrac{1+\sqrt{5}}{2})\wedge\operatorname{LOG}(1+\sqrt{2})$, we have  $\lVert w\rVert_{1}=8\log(\tfrac{1+\sqrt{5}}{2})\log(1+\sqrt{2})$ and thus
+$A_{4,2}\leq 8\log\left(\tfrac{1+\sqrt{5}}{2}\right)\log(1+\sqrt{2})\approx 3.393.$
+Hence our lower bound is no less than a third of an upper bound for $A_{4,2}$.
+The proof of the theorem is done separately for the case when  $\operatorname{Gal}(L/\mathbb{Q})$ is a Klein four-group, and the case when  $\operatorname{Gal}(L/\mathbb{Q})$ is cyclic.
 For the Klein four-group case, we define a subgroup $E$ generated by the units of the quadratic subfields of $L$.
-Using the Galois module structure of $E$, we give a lower bound of  $\lVert w\rVert_{1}$ for nonzero element  $w$ of $\bigwedge^{2}\operatorname{LOG}(E)$ and extend this result to ${\mathcal{O}_{L}^{*}}$.
+Using the Galois module structure of $E$, we give a lower bound for  $\lVert w\rVert_{1}$ for nonzero element  $w$ of $\bigwedge^{2}\operatorname{LOG}(E)$ and extend this result to $\mathcal{O}_{L}^{*}$}.
 For the cyclic group case, we use the structure of  $\mathcal{O}_{L}^{*}$} given by Hasse [5] to obtain a basis of $\bigwedge^{2}\operatorname{LOG}({\mathcal{O}_{L}^{*}})$.
 We then use the result of Pohst [8] to compute the lower bound.
-In § 2 we derive the best lower bound of  $A_{4,2}$ known so far, from the work of Costa and Friedman [3].
+In § 2 we derive the best lower bound for  $A_{4,2}$ known so far, from the work of Costa and Friedman [3].
 In the remaining sections, we prove the main theorem to give the better lower bound for $A_{4,2}$.
@@
 The work of S.S.
-was supported by College of Natural Sciences Undergraduate Internship Program from Seoul National University.
+was supported by the College of Natural Sciences Undergraduate Internship Program from Seoul National University.
 D.K.
@@
 2020R1C1C1A0100681913.
+We also thank the referee for corrections, including those concerning the cyclic case.
 2. The lower bound due to Costa and Friedman
-In this section, we derive the largest lower bound of  $A_{4,2}$ known so far, to the best of the author’s knowledge.
+In this section, we derive the largest lower bound for  $A_{4,2}$ known so far, to the best of the authors’ knowledge.
 Using Pohst’s result [8], Costa and Friedman [3] showed that for independent elements $\epsilon_{1},...,\epsilon_{j}$ of $\mathcal{O}_{L}^{*}$}, the inequality
-$\lVert\operatorname{LOG}(\epsilon_{1})\wedge\cdots\wedge\operatorname{LOG}(\epsilon_{j})\rVert_{2}>\left(\frac{n}{\gamma_{j}}\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}\right)^{j/2}$
+$\lVert\operatorname{LOG}(\epsilon_{1})\wedge\cdots\wedge\operatorname{LOG}(\epsilon_{j})\rVert_{2}>\left(\frac{n}{\gamma_{j}}\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right)\right)^{j/2}$
 holds for $1\leq j<n$, where $n=[L:\mathbb{Q}]$ and  $\gamma_{j}$ is Hermite’s constant in dimension $j$.
-If $j=n-2=\operatorname{rank}({\mathcal{O}_{L}^{*}})-1$, every element of  $\bigwedge^{n-2}\operatorname{LOG}({\mathcal{O}_{L}^{*}})$ can be written as  $d\cdot\operatorname{LOG}(\epsilon_{1})\wedge\cdots\wedge\operatorname{LOG}(\epsilon_{n-2})$ where $d\in\mathbb{Z}$ and $\operatorname{LOG}(\epsilon_{1}),...,\operatorname{LOG}(\epsilon_{n-1})$ are the basis of $\operatorname{LOG}({\mathcal{O}_{L}^{*}})$, as shown in the Lemma 28 of [2].
+If $j=n-2=\operatorname{rank}({\mathcal{O}_{L}^{*}})-1$, every element of  $\bigwedge^{n-2}\operatorname{LOG}({\mathcal{O}_{L}^{*}})$ can be written as  $d\cdot\operatorname{LOG}(\epsilon_{1})\wedge\cdots\wedge\operatorname{LOG}(\epsilon_{n-2})$ where $d\in\mathbb{Z}$ and $\operatorname{LOG}(\epsilon_{1}),...,\operatorname{LOG}(\epsilon_{n-1})$ are the basis of $\operatorname{LOG}({\mathcal{O}_{L}^{*}})$, as shown in Lemma 28 of [2].
 Therefore, for any nonzero element  $w$ of  $\bigwedge^{n-2}\operatorname{LOG}({\mathcal{O}_{L}^{*}})$, the inequality
-$\lVert w\rVert_{1}\geq\lVert w\rVert_{2}>\left(\frac{n}{\gamma_{j}}\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}\right)^{j/2}$
+$\lVert w\rVert_{1}\geq\lVert w\rVert_{2}>\left(\frac{n}{\gamma_{j}}\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right)\right)^{j/2}$
 holds. Restricting to the Galois extensions of $\mathbb{Q}$, we have
-$A_{n,n-2}>\left(\frac{n}{\gamma_{n-2}}\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}\right)^{\frac{n-2}{2}}$
+$A_{n,n-2}>\left(\frac{n}{\gamma_{n-2}}\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right)\right)^{\frac{n-2}{2}}$
 and in particular, since $\gamma_{2}=\frac{2}{\sqrt{3}}$, we obtain
-$A_{4,2}>2\sqrt{3}\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}\approx 0.802.$
+$A_{4,2}>2\sqrt{3}\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right)\approx 0.802.$
 3. The Klein four-group Case
@@
 For $i=1,2,3$, let $\operatorname{Gal}(L/\mathbb{Q}(\sqrt{d_{i}}))=\{1,\sigma_{i}\}$.
-See Figure 1.
 Let  $u_{i}>1$ be the fundamental unit of $\mathbb{Q}(\sqrt{d_{i}})$.
 Without loss of generality, let $1<u_{1}<u_{2}<u_{3}$.
-If $i$ and $j$ are distinct integers in $\{1,2,3\}$, since $\sqrt{d_{j}}\not\in\mathbb{Q}(\sqrt{d_{i}})$$\sigma_{i}\in\operatorname{Gal}(L/\mathbb{Q}(\sqrt{d_{i}}))$ sends $\sqrt{d_{j}}$ to its conjugate $-\sqrt{d_{j}}$.
+If $i$ and $j$ are distinct integers in $\{1,2,3\}$, since $\sqrt{d_{j}}\not\in\mathbb{Q}(\sqrt{d_{i}})$$\sigma_{i}\in\operatorname{Gal}(L/\mathbb{Q}(\sqrt{d_{i}}))$ sends $\sqrt{d_{j}}$ to its conjugate $\sqrt{d_{j}}$.
 Thus
 (3.1) $u_{j}\sigma_{i}(u_{j})=\operatorname{N}_{\mathbb{Q}(\sqrt{d_{j}})/\mathbb{Q}}(u_{j})=\pm 1.$
-$\mathbb{Q}$$\mathbb{Q}(\sqrt{d_{2}})$$\mathbb{Q}(\sqrt{d_{3}})$$\mathbb{Q}(\sqrt{d_{1}})$$L=\mathbb{Q}(\sqrt{d_{1}},\sqrt{d_{2}})$$\{id,\sigma_{1}\}$$\{id,\sigma_{2}\}$$\{id,\sigma_{3}\}$$\{id,\sigma_{1},\sigma_{2},\sigma_{3}\}$$\{id\}$
-Figure 1. Subfields of non-cyclic extension $L$ and corresponding Galois groups
+Let $E=\{u_{1}^{m_{1}}u_{2}^{m_{2}}u_{3}^{m_{3}}|m_{1},m_{2},m_{3}\in\mathbb{Z}\}$.
+Then we have the following lemmas, which are immediate consequences of Kubota’s classification of fundamental systems of units in real biquadratic fields [6, Satz 1].
 Lemma 3.1.
 $u_{1},u_{2},u_{3}$ are multiplicatively independent. In other words, if integers $m_{1},m_{2},m_{3}$ satisfy $u_{1}^{m_{1}}u_{2}^{m_{2}}u_{3}^{m_{3}}=1$, then $m_{1}=m_{2}=m_{3}=0$.
-Proof.
-Assume that
-$u_{1}^{m_{1}}u_{2}^{m_{2}}u_{3}^{m_{3}}=1$
-for some $m_{1},m_{2},m_{3}\in\mathbb{Z}$. Then we have
-$\sigma_{1}(u_{1})^{m_{1}}\sigma_{1}(u_{2})^{m_{2}}\sigma_{1}(u_{3})^{m_{3}}=1$
-and thus
-$(u_{1}\sigma_{1}(u_{1}))^{m_{1}}(u_{2}\sigma_{1}(u_{2}))^{m_{2}}(u_{3}\sigma_{1}(u_{3}))^{m_{3}}=1.$
-By (3.1), $u_{2}\sigma_{1}(u_{2})=\pm 1$ and $u_{3}\sigma_{1}(u_{3})=\pm 1$. Therefore
-$u_{1}^{2m_{1}}=\pm 1$
-and thus $m_{1}=0$. Similarly, $m_{2}=m_{3}=0$. Therefore, $u_{1},u_{2},u_{3}$ are multiplicatively independent. ∎
-Let $E=\{u_{1}^{m_{1}}u_{2}^{m_{2}}u_{3}^{m_{3}}|m_{1},m_{2},m_{3}\in\mathbb{Z}\}$, then we have the following lemma.
 Lemma 3.2.
 If $u\in{\mathcal{O}_{L}^{*}}$, then $u^{2}\in E$.
-Proof.
-From  $u\in{\mathcal{O}_{L}^{*}}$, we have  $\lvert N_{L/\mathbb{Q}}(u)\rvert=1$.
-The unit  $u$ of  $L$ and its conjugate  $\sigma_{1}(u)$ are units of  $L$, so  $u\sigma_{1}(u)$ is a unit of  $L$.
-Since  $u\sigma_{1}(u)=N_{L/\mathbb{Q}(\sqrt{d_{1}})}(u)\in\mathbb{Q}(\sqrt{d_{1}})$, it is also a unit of $\mathbb{Q}(\sqrt{d_{1}})$.
-Hence  $u\sigma_{1}(u)=\pm u_{1}^{m_{1}}$ for some $m_{1}\in\mathbb{Z}$.
-Similarly, $u\sigma_{2}(u)=\pm u_{2}^{m_{2}}$ and $u\sigma_{1}(u)=\pm u_{3}^{m_{3}}$ for some $m_{2},m_{3}\in\mathbb{Z}$.
-Then we have
-$u^{2}=\lvert u^{2}\cdot N_{L/\mathbb{Q}}(u)\rvert=\lvert u^{3}\sigma_{1}(u)\sigma_{2}(u)\sigma_{3}(u)\rvert=\lvert\pm u_{1}^{m_{1}}u_{2}^{m_{2}}u_{3}^{m_{3}}\rvert=u_{1}^{m_{1}}u_{2}^{m_{2}}u_{3}^{m_{3}}$
-and therefore $u^{2}\in E$. ∎
-[…]
+We first compute the 1-norm of the elements of $\bigwedge^{2}\operatorname{LOG}(E)$.
+$\operatorname{LOG}(E)$ is generated by $\operatorname{LOG}(u_{i})$’s.
+Using (3.1), we can compute $\operatorname{LOG}(u_{i})$’s in terms of $u_{i}$’s as follows:
 $\operatorname{LOG}(u_{1})=(\log(u_{1}),\ \ \log(u_{1}),-\log(u_{1}),-\log(u_{1}))$
@@
 Then the wedge product of these $\operatorname{LOG}(u_{i})$’s are:
-[…]
-[…]
-[…]
-where $X_{1}=\log(u_{2})\log(u_{3}),X_{2}=\log(u_{1})\log(u_{3}),X_{3}=\log(u_{1})\log(u_{2})$ and the order of the basis of $\bigwedge^{j}\mathbb{R}^{\mathcal{A}_{L}}$ is $\delta^{id}\wedge\delta^{\sigma_{1}},\delta^{\sigma_{2}}\wedge\delta^{\sigma_{3}},\delta^{id}\wedge\delta^{\sigma_{3}},\delta^{\sigma_{1}}\wedge\delta^{\sigma_{2}},\delta^{id}\wedge\delta^{\sigma_{2}},\delta^{\sigma_{1}}\wedge\delta^{\sigma_{3}}$.
+$\operatorname{LOG}(u_{2})\wedge\operatorname{LOG}(u_{3})=(\ \ \ \ \ 0,\ \ \ \ \ 0,\ \ 2X_{1},\ \ 2X_{1},-2X_{1},-2X_{1})$
+$\operatorname{LOG}(u_{1})\wedge\operatorname{LOG}(u_{3})=(-2X_{2},-2X_{2},\ \ 2X_{2},-2X_{2},\ \ \ \ \ 0,\ \ \ \ \ 0)$
+$\operatorname{LOG}(u_{1})\wedge\operatorname{LOG}(u_{2})=(-2X_{3},\ \ 2X_{3},\ \ \ \ \ 0,\ \ \ \ \ 0,\ \ 2X_{3},-2X_{3})$
+where $X_{1}=\log(u_{2})\log(u_{3}),X_{2}=\log(u_{1})\log(u_{3}),X_{3}=\log(u_{1})\log(u_{2})$ and the order of the basis of $\bigwedge^{2}\mathbb{R}^{\mathcal{A}_{L}}$ is $\delta^{id}\wedge\delta^{\sigma_{1}},\delta^{\sigma_{2}}\wedge\delta^{\sigma_{3}},\delta^{id}\wedge\delta^{\sigma_{3}},\delta^{\sigma_{1}}\wedge\delta^{\sigma_{2}},\delta^{id}\wedge\delta^{\sigma_{2}},\delta^{\sigma_{1}}\wedge\delta^{\sigma_{3}}$.
 Note that $X_{1}>X_{2}>X_{3}>0$.
-For $w\in\bigwedge^{2}\operatorname{LOG}(E)$$w$ can be written as
+For $\in\bigwedge^{2}\operatorname{LOG}(E)ww$ can be written as
 $w=n_{1}\cdot\operatorname{LOG}(u_{2})\wedge\operatorname{LOG}(u_{3})+n_{2}\cdot\operatorname{LOG}(u_{1})\wedge\operatorname{LOG}(u_{3})+n_{3}\cdot\operatorname{LOG}(u_{1})\wedge\operatorname{LOG}(u_{2})$
@@
 $\lVert w\rVert_{1}=$  $\,2(\lvert n_{2}X_{2}+n_{3}X_{3}\rvert+\lvert n_{2}X_{2}-n_{3}X_{3}\rvert+\lvert n_{1}X_{1}+n_{2}X_{2}\rvert+\lvert n_{1}X_{1}-n_{2}X_{2}\rvert$
-$+\lvert n_{1}X_{1}+n_{3}X_{3}\rvert+\lvert n_{1}X_{1}-n_{3}X_{3}\rvert)$
-$=$  $\,4(\max\{\lvert n_{2}X_{2}\rvert,\lvert n_{3}X_{3}\rvert\}+\max\{\lvert n_{1}X_{1}\rvert,\lvert n_{2}X_{2}\rvert\}+\max\{\lvert n_{1}X_{1}\rvert,\lvert n_{3}X_{3}\rvert\})$
+[…]
+[…]
 where the second equality is from the following equation.
@@
 holds. Similarly, the above inequality also holds when the right hand side is $8\lvert n_{2}X_{2}\rvert$ or $8\lvert n_{3}X_{3}\rvert$. Therefore
-$\lVert w\rVert_{1}$ $=4(\max\{\lvert n_{2}X_{2}\rvert,\lvert n_{3}X_{3}\rvert\}+\max\{\lvert n_{1}X_{1}\rvert,\lvert n_{2}X_{2}\rvert\}+\max\{\lvert n_{1}X_{1}\rvert,\lvert n_{3}X_{3}\rvert\})$
+[…]
 $\geq 8\max\{\lvert n_{1}X_{1}\rvert,\lvert n_{2}X_{2}\rvert,\lvert n_{3}X_{3}\rvert\}$
@@
 Hence we have
-(3.4) $\lVert w\rVert_{1}=\frac{1}{2}\lVert 2w\rVert_{1}\geq 4\log\left(\frac{1+\sqrt{5}}{2}\right)\log(1+\sqrt{2})\approx 1.697.$
+(3.4) $\lVert w\rVert_{1}=\frac{1}{2}\lVert 2w\rVert_{1}\geq 4\log\left(\tfrac{1+\sqrt{5}}{2}\right)\log(1+\sqrt{2})\approx 1.697.$
 For the case $L=\mathbb{Q}(\sqrt{5},\sqrt{13})$$u_{1}=\frac{1+\sqrt{5}}{2},u_{2}=\frac{3+\sqrt{13}}{2},u_{3}=8+\sqrt{65}$.
 Similarly, $u_{1},u_{2},\sqrt{u_{1}u_{2}u_{3}}$ generate  $\mathcal{O}_{L}^{*}$} and therefore
-(3.5) $\lVert w\rVert_{1}=\frac{1}{2}\lVert 2w\rVert_{1}\geq 4\log\left(\frac{1+\sqrt{5}}{2}\right)\log(\frac{3+\sqrt{13}}{2})\approx 2.300$
+(3.5) $\lVert w\rVert_{1}=\frac{1}{2}\lVert 2w\rVert_{1}\geq 4\log\left(\tfrac{1+\sqrt{5}}{2}\right)\log\left(\tfrac{3+\sqrt{13}}{2}\right)\approx 2.300$
 for any nonzero element $w$ of $\bigwedge^{2}\operatorname{LOG}({\mathcal{O}_{L}^{*}})$.
@@
 by (3.3).
-Excluding $\mathbb{Q}(\sqrt{2},\sqrt{5})$ and $\mathbb{Q}(\sqrt{5},\sqrt{13})$, $\mathbb{Q}(\sqrt{3},\sqrt{5})$ has the smallest possible value of  $\log(u_{1})\log(u_{2})$, which is  $\log\left(\frac{1+\sqrt{5}}{2}\right)\log(2+\sqrt{3})$.
-This is from the Lemma 3.3 below and the fact that $\log\left(\frac{1+\sqrt{5}}{2}\right)\log(2+\sqrt{3})<\log(1+\sqrt{2})\log(\frac{3+\sqrt{13}}{2})$.
+Excluding $\mathbb{Q}(\sqrt{2},\sqrt{5})$ and $\mathbb{Q}(\sqrt{5},\sqrt{13})$, $\mathbb{Q}(\sqrt{3},\sqrt{5})$ has the smallest possible value of  $\log(u_{1})\log(u_{2})$, which is  $\log\left(\tfrac{1+\sqrt{5}}{2}\right)\log(2+\sqrt{3})$.
+This is from the Lemma 3.3 below and the fact that $\log\left(\tfrac{1+\sqrt{5}}{2}\right)\log(2+\sqrt{3})<\log(1+\sqrt{2})\log\left(\frac{3+\sqrt{13}}{2}\right)$.
 Hence from (3.6), we see that
-$\lVert w\rVert_{1}\geq 2\log\left(\frac{1+\sqrt{5}}{2}\right)\log(2+\sqrt{3})\approx 1.267.$
+$\lVert w\rVert_{1}\geq 2\log\left(\tfrac{1+\sqrt{5}}{2}\right)\log(2+\sqrt{3})\approx 1.267.$
 With (3.4) and (3.5), we conclude that if $\operatorname{Gal}(L/\mathbb{Q})$ is Klein four-group,
-$\lVert w\rVert_{1}\geq 2\log\left(\frac{1+\sqrt{5}}{2}\right)\log(2+\sqrt{3})>3\sqrt{3}\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}$
+$\lVert w\rVert_{1}\geq 2\log\left(\tfrac{1+\sqrt{5}}{2}\right)\log(2+\sqrt{3})>2\sqrt{6}\log\left(\tfrac{1+\sqrt{5}}{2}\right)^{2}$
 for any nonzero $w\in\bigwedge^{2}\operatorname{LOG}({\mathcal{O}_{L}^{*}})$.
@@
 For square-free integer $m$, let $v_{m}>1$ be the fundamental unit of $\mathbb{Q}(\sqrt{m})$.
-Then  $\frac{1+\sqrt{5}}{2},1+\sqrt{2},\frac{3+\sqrt{13}}{2},2+\sqrt{3}$ are the four smallest possible values of $v_{m}$.
+Then  $\tfrac{1+\sqrt{5}}{2},1+\sqrt{2},\tfrac{3+\sqrt{13}}{2},2+\sqrt{3}$ are the four smallest possible values of $v_{m}$.
 Proof.
@@
 Let $l$ be the unique quadratic subfield of $L$.
-See Figure 2.
 Let $u_{l}>1$ be the fundamental unit of $l$.
@@
 First we define relative units as in [5].
-$\mathbb{Q}$$l$$L$$\{1,\sigma^{2}\}$$G=\{1,\sigma,\sigma^{2},\sigma^{3}\}$$\{1\}$
-Figure 2. Subfields of cyclic extension $L$ and corresponding Galois groups
 Definition 4.1.
@@
 Let $Q=[{\mathcal{O}_{L}^{*}}:E^{L}]$.
-The following lemmas are from Hasse [5].
+The following two lemmas are due to Hasse [5, Satz 16 and 22].
 Lemma 4.2.
 The index $Q$ is either 1 or 2. Furthermore,  $Q=2$ if and only if there exists $u_{*}\in{\mathcal{O}_{L}^{*}}$ such that $N_{L/l}(u_{*})=\pm u_{l}$.
-Proof.
-See Satz 16, p.37 of Hasse [5]. ∎
 Lemma 4.3.
-There exists an element $u_{0}$ of $E_{L/l}$ such that $-1,u_{0}$ and its conjugate $\sigma(u_{0})$ generates $E_{L/l}$.
-Proof.
-See Satz 22, p.40 of Hasse [5]. ∎
+There exists an element $u_{0}$ of $E_{L/l}$ such that $1,u_{0}$ and its conjugate $\sigma(u_{0})$ generates $E_{L/l}$.
 Define $u_{0}$ as in the lemma. We have
@@
 (4.1) $\begin{split}\operatorname{LOG}(u_{l})=(W_{1},-W_{1},W_{1},-W_{1})\\ \operatorname{LOG}(u_{0})=(W_{2},W_{3},-W_{2},-W_{3})\\ \operatorname{LOG}(\sigma(u_{0}))=(W_{3},-W_{2},-W_{3},W_{2})\end{split}$
-where $W_{1}=\log(u_{l}),W_{2}=\log(\lvert u_{0}\rvert),W_{3}=\log(\lvert\sigma(u_{0})\rvert)$.
-Here, the order of the basis of $\mathbb{R}^{\mathcal{A}_{L}}$ is $\delta^{id},\delta^{\sigma},\delta^{\sigma^{2}},\delta^{\sigma^{3}}$.
-Note that $W_{1},W_{2},W_{3}$ are nonzero since $u_{l},u_{0},\sigma(u_{0})$ are not ±$\pm 1$.
-Then the wedge product of these terms are:
+[…]
 $\operatorname{LOG}(u_{l})\wedge\operatorname{LOG}(u_{0})=$ $(\ \ Y_{4},-Y_{4},Y_{5},\ \ Y_{5},-Y_{2},\ \ Y_{3})$
 $\operatorname{LOG}(u_{l})\wedge\operatorname{LOG}(\sigma(u_{0}))=$ $(-Y_{5},\ \ Y_{5},Y_{4},\ \ Y_{4},-Y_{3},-Y_{2})$
-[…]
+$\operatorname{LOG}(u_{0})\wedge\operatorname{LOG}(\sigma(u_{0}))=$ $(-Y_{1},-Y_{1},Y_{1},-Y_{1},\ \ \ \ 0,\ \ \ \ 0)$
 where $Y_{1}=W_{2}^{2}+W_{3}^{2},Y_{2}=2W_{1}W_{2},Y_{3}=2W_{1}W_{3},Y_{4}=W_{1}W_{2}+W_{1}W_{3},Y_{5}=W_{1}W_{2}-W_{1}W_{3}$.
@@
 In particular, we have
-(4.2) $f(n_{1},n_{2},n_{3})\geq 4\lvert n_{3}Y_{3}\rvert=4\lvert n_{3}\rvert(W_{2}^{2}+W_{3}^{2})$
+(4.2) $f(n_{1},n_{2},n_{3})\geq 4\lvert n_{3}Y_{1}\rvert=4\lvert n_{3}\rvert(W_{2}^{2}+W_{3}^{2})$
 and
@@
 and applying (3.2) once more, we have
-(4.3) $f(n_{1},n_{2},n_{3})\geq 2W_{1}(2\max\{\lvert W_{2}\rvert,\lvert W_{3}\rvert\}+\lvert W_{2}\rvert+\lvert W_{3}\rvert)$
-Using these results, we prove the Theorem 1.2 for each case $Q=1$ and $Q=2$.
+[…]
+Using these results, we prove Theorem 1.2 for each case $Q=1$ and $Q=2$.
 Lemma 4.4.
-Let $X,Y\in\mathbb{R}$. For any $(m,n)\in\mathbb{Z}^{2}\backslash\{(0,0)\}$, the following inequality
-$\lvert mX+nY\rvert+\lvert nX-mY\rvert\geq\lvert X\rvert+\lvert Y\rvert$
-holds.
+[…]
+$\lvert mX+nY\rvert+\lvert nX-mY\rvert\geq\lvert X\rvert+\lvert Y\rvert$.
 Proof.
 The inequality is trivial if $X=0$ or $Y=0$. Assume that they are both nonzero. We may further assume that they are positive and $X\geq Y$. Let $Z:=Y/X\leq 1$. Since
-$\lvert mX+nY\rvert+\lvert nX-mY\rvert=\lvert X\rvert(\lvert m+nZ\rvert+\lvert n-mZ\rvert),$
+[…]
 we only need to show the inequality
 $\lvert m+nZ\rvert+\lvert n-mZ\rvert\geq 1+Z.$
-Without loss of generality, assume $m\geq 0$. If $m=0$$\lvert nZ\rvert+\lvert n\rvert\geq 1+Z$ for any nonzero integer $n$. If $m\geq 2$,
-$\lvert m+nZ\rvert+\lvert n-mZ\rvert$ $\geq\lvert m+nZ\rvert+\lvert Z(n-mZ)\rvert$
-$\geq\lvert(m+nZ)-Z(n-mZ)\rvert$
-$=\lvert m(1+Z^{2})\rvert$
-$\geq m$
-$\geq 2$
-$\geq 1+Z.$
+[…]
+[…]
+$=\lvert m(1+Z^{2})\rvert\;\geq m\;\geq 2\;\geq 1+Z.$
 Now consider the case $m=1$. The case $n=0$ is trivial. If $n\leq-1$, then
-$\lvert 1+nZ\rvert+\lvert n-Z\rvert\geq{n-Z}\geq 1+Z$
+$\lvert 1+nZ\rvert+\lvert n-Z\rvert\geq{-n+Z}\geq 1+Z,$
 and if $n\geq 1$,
@@
 and if $n_{3}=0$, then $(n_{1},n_{2})\neq(0,0)$, so (4.3) gives
-$\lVert w\rVert_{1}\geq 2W_{1}(2\max\{\lvert W_{2}\rvert,\lvert W_{3}\rvert\}+\lvert W_{2}\rvert+\lvert W_{3}\rvert).$
+[…]
 Therefore we have
-(4.4) $\lVert w\rVert_{1}\geq\min\{4(W_{2}^{2}+W_{3}^{2}),2W_{1}(2\max\{\lvert W_{2}\rvert,\lvert W_{3}\rvert\}+\lvert W_{2}\rvert+\lvert W_{3}\rvert)\}$
+[…]
 for any nonzero $w\in\bigwedge\operatorname{LOG}({\mathcal{O}_{L}^{*}})$.
@@
 Theorem 4.5.
-Let $L$ be totally real number field and $u\in{\mathcal{O}_{L}^{*}}$ with $u\neq\pm 1$. Then
-$\lVert\operatorname{LOG}(u)\rVert_{2}^{2}\geq[L:\mathbb{Q}]\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}$
-holds.
+Let $L$ be a totally real number field and $u\in{\mathcal{O}_{L}^{*}}$ with $u\neq\pm 1$. Then we have
+$\lVert\operatorname{LOG}(u)\rVert_{2}^{2}\geq[L:\mathbb{Q}]\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right).$
 Proof.
-Bemerkung in p.102 of Pohst [7] states that the Lemma in p.98 applies for  $M$ of the eqaution (10) in p.97. Then the inequality (11) of the Lemma gives the required result. ∎
+Bemerkung in p. 102 of Pohst [8] states that the Lemma in p. 98 applies for  $M$ of the equation (10) in p. 97. Then the inequality (11) of the Lemma gives the required result. ∎
 Then from the above theorem and (4.1),
-(4.5) $W_{2}^{2}+W_{3}^{2}=\frac{1}{2}\lVert\operatorname{LOG}(u_{0})\rVert_{2}^{2}\geq 2\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}.$
-Under this constraint and the fact that $W_{2},W_{3}$ are nonzero, an elementary consideration yields
-$2\max\{\lvert W_{2}\rvert,\lvert W_{3}\rvert\}+\lvert W_{2}\rvert+\lvert W_{3}\rvert>3\sqrt{2}\log\left(\frac{1+\sqrt{5}}{2}\right).$
+(4.5) $W_{2}^{2}+W_{3}^{2}=\frac{1}{2}\lVert\operatorname{LOG}(u_{0})\rVert_{2}^{2}\geq 2\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right).$
+Under this constraint, the following elementary fact will be useful.
+Lemma 4.6.
+Let $x,y\in\mathbb{R}$ with $x^{2}+y^{2}\geq R^{2}$ for some positive real number $R$. Then,
+$2\max\{\lvert x\rvert,\lvert y\rvert\}+\lvert x\rvert+\lvert y\rvert\geq 2\sqrt{2}R.$
+Proof.
+Without loss of generality, assume $x\geq y\geq 0$ and $x^{2}+y^{2}=R^{2}$. Then there exists $\theta\in[0,\tfrac{\pi}{4}]$ such that $x=R\cos\theta$ and $y=R\sin\theta$. Hence
+$2\max\{\lvert x\rvert,\lvert y\rvert\}+\lvert x\rvert+\lvert y\rvert=3x+y=\sqrt{10}R\sin(\theta+\phi),$
+where $\phi=\arctan(3)$.
+Since $\theta+\phi$ ranges over $\arctan(3),\arctan(3)+\tfrac{\pi}{4}]\subset(0,\pi)$ and $\sin(\phi)=\tfrac{3}{\sqrt{10}}$ is larger than $\sin(\phi+\tfrac{\pi}{4})=\tfrac{2}{\sqrt{5}}$, the minimum of $\sin(\theta+\phi)$ occurs at $\theta=\tfrac{\pi}{4}$.
+The minimum value is therefore $\sqrt{10}\,R\cdot\tfrac{2}{\sqrt{5}}\;=\;2\sqrt{2}\,R$.
+∎
+Combining Lemma 4.6 with (4.5), we obtain
+$2\max\{\lvert W_{2}\rvert,\lvert W_{3}\rvert\}+\lvert W_{2}\rvert+\lvert W_{3}\rvert\geq 4\log\left(\tfrac{1+\sqrt{5}}{2}\right).$
 Since $u_{l}$ is a fundamental unit of quadratic field, by Lemma 3.3,
-(4.6) $W_{1}=\log(u_{l})\geq\log\left(\frac{1+\sqrt{5}}{2}\right)$
-holds, and with previous inequality,
-$2W_{1}(2\max\{\lvert W_{2}\rvert,\lvert W_{3}\rvert\}+\lvert W_{2}\rvert+\lvert W_{3}\rvert)>6\sqrt{2}\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}.$
+(4.6) $W_{1}=\log(u_{l})\geq\log\left(\tfrac{1+\sqrt{5}}{2}\right)$
+holds, and with the previous inequality,
+$2W_{1}(2\max\{\lvert W_{2}\rvert,\lvert W_{3}\rvert\}+\lvert W_{2}\rvert+\lvert W_{3}\rvert)\geq 8\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right).$
 With (4.5), the inequality (4.4) becomes
-(4.7) $\lVert w\rVert_{1}\geq\min\left\{8\log\left(\frac{1+\sqrt{5}}{2}\right)^{2},6\sqrt{2}\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}\right\}=8\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}$
+(4.7) $\lVert w\rVert_{1}\geq 8\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right)$
 for any nonzero $w\in\bigwedge^{2}\operatorname{LOG}({\mathcal{O}_{L}^{*}})$. Thus Theorem 1.2 holds for the case $Q=1$.
 4.2. The Case Q=2
-We use the following lemma from Hasse [5].
+We use the following lemma due to Hasse [5, Satz 28].
 Lemma 4.7.
-If $Q=2$, there exists a unique unit $u_{*}>0$ of $L$ that satisfy
-(4.8) $N_{L/l}(u_{*})=u_{*}\sigma^{2}(u_{*})=\pm u_{l},\quad u_{*}\sigma(u_{*})=\pm u_{0}.$
-Proof.
-See Satz 28, p.45 of Hasse [5]. ∎
+If $Q=2$, there exists a unique unit $u_{*}>0$ of $L$ such that
+$|N_{L/l}(u_{*})|=u_{l},\qquad|u_{*}\sigma(u_{*})|=\lvert u_{0}\rvert$.
 Define $u_{*}$ as in the lemma. Then
@@
 and therefore  $\sigma(u_{0})=\pm u_{l}u_{0}u_{*}^{-2}$.
-The proof of Lemma 4.2 shows that when $Q=2$, every unit $u\in{\mathcal{O}_{L}^{*}}$ can be written uniquely as the form ±$\pm u_{*}^{n_{0}}u_{l}^{n_{1}}u_{0}^{n_{2}}\sigma(u_{0})^{n_{3}}$ where $n_{0}$ is 00 or $1$ and $n_{1},n_{2},n_{3}$ are integers.
-Thus  ${\mathcal{O}_{L}^{*}}$ is generated by $-1,u_{l},u_{0},u_{*}$.
+The proof of Lemma 4.2 shows that when $Q=2$, every unit $u\in{\mathcal{O}_{L}^{*}}$ can be written uniquely as the form $\pm u_{*}^{n_{0}}u_{l}^{n_{1}}u_{0}^{n_{2}}\sigma(u_{0})^{n_{3}}$ where $n_{0}$ is 0 or $1$ and $n_{1},n_{2},n_{3}$ are integers.
+Thus  $\mathcal{O}_{L}^{*}$} is generated by $1,u_{l},u_{0},u_{*}$.
 Hence  $\bigwedge^{2}\operatorname{LOG}({\mathcal{O}_{L}^{*}})$ is generated by $\operatorname{LOG}(u_{l})\wedge\operatorname{LOG}(u_{0}),\operatorname{LOG}(u_{l})\wedge\operatorname{LOG}(u_{*})$ and  $\operatorname{LOG}(u_{0})\wedge\operatorname{LOG}(u_{*})$.
@@
 We now proceed to give a lower bound for $f(n_{1},n_{2},n_{3})$ where $n_{1}+n_{2}+n_{3}$ is even and $(n_{1},n_{2},n_{3})\neq(0,0,0)$.
-If $(n_{1},n_{2})=0$, then $\lvert n_{3}\rvert\geq 2$, so (4.2) gives
+If $(n_{1},n_{2})=(0,0)$, then $\lvert n_{3}\rvert\geq 2$, so (4.2) gives
 $f(n_{1},n_{2},n_{3})\geq 8(W_{2}^{2}+W_{3}^{2}).$
 Since (4.5) still holds in the $Q=2$ case, we have
-(4.12) $f(n_{1},n_{2},n_{3})\geq 8\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}.$
+(4.11) $f(n_{1},n_{2},n_{3})\geq 16\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right).$
 Note that (4.6) also still holds.
-Now consider the case $(n_{1},n_{2})\neq(0,0)$. Since $\operatorname{LOG}(u_{l}),\operatorname{LOG}(u_{0}),\operatorname{LOG}(\sigma(u_{0}))~{}$ are perpendicular as seen in (4.1), we have
+Now consider the case $(n_{1},n_{2})\neq(0,0)$.
+Since $\operatorname{LOG}(u_{l}),\operatorname{LOG}(u_{0}),\operatorname{LOG}(\sigma(u_{0}))$~ are perpendicular as follows from (4.1), we have from (4.9)
 $\lVert\operatorname{LOG}(u_{*})\rVert_{2}^{2}$ $=\frac{1}{4}(\lVert\operatorname{LOG}(u_{l})\rVert_{2}^{2}+\lVert\operatorname{LOG}(u_{0})\rVert_{2}^{2}+\lVert\operatorname{LOG}(\sigma(u_{0}))\rVert_{2}^{2})$
-$=W_{1}^{2}+W_{2}^{2}+W_{3}^{2}$
-from (4.10). Then Theorem 4.5 gives
-$W_{1}^{2}+W_{2}^{2}+W_{3}^{2}\geq 4\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}.$
-Under the constraint of above inequality, (4.5), (4.6), and the fact that  $W_{1},W_{2},W_{3}$ are nonzero, an elementary consideration yields
-$2W_{1}(2\max\{\lvert W_{2}\rvert,\lvert W_{3}\rvert\}+\lvert W_{2}\rvert+\lvert W_{3}\rvert)>6\sqrt{3}\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}.$
+$=W_{1}^{2}+W_{2}^{2}+W_{3}^{2}.$
+Then Theorem 4.5 gives
+$W_{1}^{2}+W_{2}^{2}+W_{3}^{2}\geq 4\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right).$
+Letting $S=\sqrt{W_{2}^{2}+W_{3}^{2}}$, the inequality becomes
+(4.12) $W_{1}^{2}+S^{2}\geq 4\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right),$
+while (4.5) gives $S\geq\sqrt{2}\log\left(\tfrac{1+\sqrt{5}}{2}\right)$.
+We claim that
+(4.13) $W_{1}S\geq\sqrt{3}\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right).$
+Indeed, if $W_{1}\geq\sqrt{2}\log\left(\tfrac{1+\sqrt{5}}{2}\right)$, then since $S\geq\sqrt{2}\log\left(\tfrac{1+\sqrt{5}}{2}\right)$,
+$W_{1}S\geq 2\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right)>\sqrt{3}\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right).$
+Otherwise, by(4.6), we have $W_{1}\in\left[\log\left(\tfrac{1+\sqrt{5}}{2}\right),\sqrt{2}\log\left(\tfrac{1+\sqrt{5}}{2}\right)\right)$. Then
+$W_{1}S\geq\sqrt{W_{1}^{2}(4\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right)-W_{1}^{2})}=\sqrt{4\log^{4}\left(\tfrac{1+\sqrt{5}}{2}\right)-\left(W_{1}^{2}-2\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right)\right)^{2}}.$
+This expression attains its minimum at $W_{1}=\log\left(\tfrac{1+\sqrt{5}}{2}\right)$, yielding $\sqrt{3}\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right)$. This proves the claim.
+Combining Lemma  4.6 with (4.13), we obtain
+$2W_{1}(2\max\{\lvert W_{2}\rvert,\lvert W_{3}\rvert\}+\lvert W_{2}\rvert+\lvert W_{3}\rvert)\geq 4\sqrt{2}W_{1}S\geq 4\sqrt{6}\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right).$
 With (4.3), we have
-$f(n_{1},n_{2},n_{3})>6\sqrt{3}\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}.$
-Since $8\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}>6\sqrt{3}\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}$, by (4.12), the above inequality holds for any $(n_{1},n_{2},n_{3})\neq(0,0,0)$ with $n_{1}+n_{2}+n_{3}$ even.
+$f(n_{1},n_{2},n_{3})\geq 4\sqrt{6}\log\left(\tfrac{1+\sqrt{5}}{2}\right)^{2}.$
+Since $16\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right)>4\sqrt{6}\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right)$, by (4.11), the above inequality holds for any $(n_{1},n_{2},n_{3})\neq(0,0,0)$ with $n_{1}+n_{2}+n_{3}$ even.
 Therefore, with (4.10),
-$\lVert w\rVert_{1}>3\sqrt{3}\log\left(\frac{1+\sqrt{5}}{2}\right)^{2}$
+$\lVert w\rVert_{1}\geq 2\sqrt{6}\log^{2}\left(\tfrac{1+\sqrt{5}}{2}\right)$
 for any nonzero $w\in\bigwedge^{2}\operatorname{LOG}({\mathcal{O}_{L}^{*}})$. This concludes the proof of Theorem 1.2.